THE SPECTROSCOPY AND AB INITIO CALCULATION OF THE C_{3} − AR COMPLEX

GUIQIU ZHANG, A. MEBEL, AND \underline{YEN-CHU HSU}, INSTITUTE OF ATOMIC AND MOLECU-LAR SCIENCES, ACADEMIA SINICA, P. O. BOX 23-166, TAIPEI 107, TAIWAN, R. O. C..

The C_{3}Ar van der Waals complex has been produced by photolyzing propargyl bromide at 193 nm in a supersonic molecular beam and probed by the laser-induced fluorescence. The spectroscopy of the ground electronic state of this complex was obtained by the wavelength-resolved emission from the Ã state. In this preliminary report, the observed vibrational levels of the C_{3}Ar complex were assigned based upon our ab initio results calculated at the MP2/aug-cc-pvtz level.